OpenStudy (anonymous):

Convert 33/4 to a percent

OpenStudy (anonymous):

help I dont under stand math

OpenStudy (anonymous):

for such sums multiply the fraction by 100 and put percent sign after the result...

OpenStudy (anonymous):

\[\frac{ 33 }{ 4 }*100=\frac{ 3300 }{ 4 }=825%\]

OpenStudy (anonymous):

I still dont understand what you mean.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

so answer is 825%

OpenStudy (anonymous):

Harkirat Thank you

OpenStudy (anonymous):

U r welcome...☺

OpenStudy (anonymous):

Harkirat can you walk me through this please. Solve the following: 56 is ______% of 125

OpenStudy (anonymous):

some one help me please

OpenStudy (anonymous):

let 56 be x% of 125 This can be written in mathematical terms as \[\frac{ x }{ 100 }*125=56\]

OpenStudy (anonymous):

Now to find the value we have to isolate it by removing the numbers attached to it. we do this as follows...